30. The pH of a solution which is 0.01M HC6H706 and 0.1M NaC6H706 is: A. 5.10- B. 8.90€ C. 4.10

please explain it...
Answer #1

The given acid is actually ascorbic acid having pKa value = 4.10.

Since a weak acid and its salt together makes a buffer solution, so, we have to use Henderson-Hasselbalch equation for getting the pH.

The HH equation for acidic buffer is:

salt pH = pka +log acid

Here, [salt] = 0.1 M and [acid]= 0.01 M

0.1 Thus, pH = 4.10+ logº = 4.10 + log10 = 4.10+1 = 5.10

So, correct answer is A.